• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name riordan

    Log entry time 17:55:54 on October 08, 2010

    Entry number 333947

    keyword=Plans for Tonight

    The NMR for the HRS has been restored and we are ready to run the
    spectrometers.  We are presently waiting for Hall C to finish cooling
    their target so we can start cooling down ours.  I've been informed by
    the acting hall leader that Hall C is deemed to have cryo priority, so we
    must wait for it to become available.
    
    JP as agreed to cool down the target if it is ready before midnight. 
    Otherwise, we will have to ask Dave Meekins to come in and do it on the
    weekend.
    
    If we do not cool down, we will leave the hall open for tonight. 
    Otherwise, we will execute a test plan for the HRS.  Swing shift, owl
    shift, and day shift tomorrow is on standby.